OpenStudy (anonymous):

i typed in a problem and i got this answer can somebody explain this to me PLEASE ?? 2.56e+22

OpenStudy (anonymous):

i believe i can help you, but i need to know the original problem

OpenStudy (anonymous):

(6.25*10^-12)^-2

OpenStudy (anonymous):

so this is 6.25 times 10 raised to the -12 power inside the parentheses

OpenStudy (anonymous):

then raise that result to the -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. so how did you proceed?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

by proceed you mean start the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok wow i feel stupid ok

OpenStudy (anonymous):

ok first you distribute the -2

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

i am so sorry some ones at my door one minute please

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

ok they left im so sorry

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

please, dont feel uncomfortable. you are fine

OpenStudy (anonymous):

which will be 6.25^-2 times 10^24

OpenStudy (anonymous):

sure. that looks good so far

OpenStudy (anonymous):

0.0256 times....

OpenStudy (anonymous):

ok it says 1e+24

OpenStudy (anonymous):

when i tried to type in 10^24

OpenStudy (anonymous):

1e2+24 means 10^24

OpenStudy (anonymous):

so e2 is the same as ^?

OpenStudy (anonymous):

ooops. 1e+24=10^24

OpenStudy (anonymous):

ok so e is ^

OpenStudy (anonymous):

it is simply notation

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

raised to the exponent of 24

OpenStudy (anonymous):

ok well thats less confusing

OpenStudy (anonymous):

it seems like your math skills and order of operations are in order. this was just a notation issue.

OpenStudy (anonymous):

ok its just im taking algebra 1 as a highschool credit and its the 2nd week and its sooooo confusing

OpenStudy (anonymous):

now the 6.25 raised to the -2 gave you the .0256

OpenStudy (phi):

on calculators or in programming languages e is used to mean 10 to the exponent so 2e3 means 2 times 10^3 or 2000 and 2e-3 means 2 times 10^-3 or 2 times 1/1000 or 1/500

OpenStudy (anonymous):

thanks phi

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

i don t think we are quite through yet

OpenStudy (anonymous):

i was talking to phi

OpenStudy (anonymous):

so back to the problem

OpenStudy (anonymous):

to get the answer, there is a final step

OpenStudy (anonymous):

.0256=2.56x10^-2

OpenStudy (anonymous):

so thats it?

OpenStudy (anonymous):

and we are multiplying this times 10^24

OpenStudy (anonymous):

it says to simplify and write it in scientific notation

OpenStudy (anonymous):

so now we add the exponents of 10

OpenStudy (anonymous):

by the laws of exponents, when you multiply two exponents with the same base, in this case 10, you add the exponents

OpenStudy (anonymous):

add them with what theres no more exponents left

OpenStudy (anonymous):

remember that i wrote for you .0256=2.56x10^-2;

OpenStudy (anonymous):

now multiply that time 1x10^24

OpenStudy (anonymous):

10^24

OpenStudy (anonymous):

yes, 1x10^24=10^24

OpenStudy (anonymous):

i wrote it that way so that you can see that the 2.56 has 1 to multiply by

OpenStudy (anonymous):

sooo its 2.56x10^24

OpenStudy (anonymous):

so the product looks like this; 1x2.56x10^24x10^-2

OpenStudy (anonymous):

not quite, you have to account for the 10^-2

OpenStudy (anonymous):

?!?

OpenStudy (anonymous):

look at the above product.. dont give up... you are nearly there

OpenStudy (anonymous):

iam COMPLETLY lost

OpenStudy (anonymous):

add the exponents of the 10's

OpenStudy (anonymous):

22

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

.0256=2.56x10^-2 and 1x10^24=10^24; then we take their product;

OpenStudy (anonymous):

(2.56x10^-2)(1x10^24); now use the commutative property to rearrange the product like so; (2.56x1)(10^-2x10^24). Does this make sense to you?

OpenStudy (anonymous):

ok my moms home now she can can help me i just cant do this on the computer it to confusing

OpenStudy (anonymous):

i"am sooooo sorry

OpenStudy (anonymous):

thank you for your time and help i just have to see it worked out on paper

OpenStudy (anonymous):

take care.

OpenStudy (phi):

maybe this video and the next one after it will help http://www.khanacademy.org/math/algebra/exponents-radicals/v/scientific-notation-i

OpenStudy (phi):

and this one (bring lots of popcorn) http://www.khanacademy.org/math/algebra/exponents-radicals/v/scientific-notation-examples